OpenStudy (anonymous):

Solve 7 cos(2 x) = 5 for the smallest three positive solutions.

OpenStudy (anonymous):

\(\Large 7 cos(2 x) = 5 \Rightarrow cos(2x)=\frac{5}{7}=cos(\alpha)\) But, I can't find \(\Large \alpha\)